You are launching water balloons at a rival team using a large slingshot. The other team is set up on the opposite side of a flat‐topped building that is 30.0 ft tall and 50.0 ft wide. Your reconnaissance team has reported that the opposition is set up 10.0 m from the wall of the building. Your balloon launcher is calibrated for launch speeds that can reach as high as 115 mph at angles between 0 and 85.0° from the horizontal. Since a direct shot is not possible (the opposing team is on the opposite side of the building), you plan to splash the other team by making a balloon explode on the ground near them.  If your launcher is located 55.0 m from the building (opposite side as the opposing team), what should your launch velocity be ((a) magnitude and (b) direction) to land a balloon 5.0 meters beyond the opposing team with maximum impact (i.e. maximum vertical speed)? 


1
Expert's answer
2021-01-28T14:01:38-0500
"v_{max}=115(0.44704)=51.41\frac{m}{s}"

"R=55+15.24+10+5=85.24\ m"

"R=\frac{v_0^2\sin{2\theta}}{g}\\85.24=\frac{51.41^2}{9.8}\sin{2\theta}\\\theta=0.5\arcsin{\frac{(85.24)9.8}{51.41^2}}=80.8\degree"